# RAMAC: Multimodal Risk-Aware Offline Reinforcement Learning and the Role of Behavior Regularization

### v1

@@ 4.1 Prior-Anchored Perturbation and its limitations
where $\zeta_{\psi}$ is a learned residual (optimized to increase $Q$ or CVaR) and the norm bound $\Phi$ keeps updates closed to the anchor. Define the anchor support $\mathcal{S}_{G}(s)$ (the region in action space where $G_{\phi}(\cdot\mid s)$ places mass), the full action space $\mathbb{R}^{d}$, and the $\Phi$-radius ball of $b$
@@
where $\mathrm{dist}(x,A)\!\mathrel{\mathop{:}}=\!\inf_{y\in A}\|x-y\|_{2}$ denotes Euclidean distance. OOD can still occur when this margin fails. This method provides a convenient local improvement rule, however, prior work has observed it suffers from poor mode coverage in multimodal action spaces (Wang et al., 2023). In addition to the identified limitations, we show distinct geometric weakness that can occur even without multimodality; having multiple modes merely magnifies the effect.
  • “Thin” support near $b$: “Thin” means the local supported region around $b$ is narrow; formally, the margin $m(b)\mathrel{\mathop{:}}=\mathrm{dist}\!\big(b,\mathbb{R}^{d}\!\setminus\!\mathcal{S}_{G}(s)\big)$ is small. If $m(b)\leq\Phi$, the ball $B_{\Phi}(b)$ overlaps the outside of $\mathcal{S}_{G}(s)$, so some $a=b+\zeta_{\psi}$ become OOD even though $\|\zeta_{\psi}\|\leq\Phi$.
  • Nonconvex support: “Nonconvex” means $\mathcal{S}_{G}(s)$ is not a convex set (e.g., a ring/annulus with a hole). Even if $b\in\mathcal{S}_{G}(s)$, a ball around $b$ may jump outside through a nearby concavity or hole whenever the margin $m(b)$ is small.
  • Gradient pushes off the data surface: The residual $\zeta_{\psi}(s,b)$ is trained to increase $Q$ or CVaR and is not constrained to be tangent to the data manifold. Consequently, gradients can point along the manifold’s normal direction, driving $a=b+\zeta_{\psi}$ toward the ball boundary ($\|\zeta_{\psi}\|\!\approx\!\Phi$). If $m(b)\!\leq\!\Phi$, these updates cross into OOD region.
@@ 4.2 Behavior‑Regularized Objective: Why it works better
This shows that shrinking forward KL via BC can suppress per–state OOD, controlled by the selection of $\eta$ (Eq. 11), hence, avoiding an important challenge of offline RL discussed in Sec. 2 (Proof appears in App. A).
We can extend this observation to provide insight on one possible mechanism through which more expressive policies can support better performance in offline-RL settings. Consider the multi-modal data $D^{\text{M}}$, uni-modal policy, $\pi_{\omega}^{\text{U}}$, and multi-modal policy, $\pi_{\xi}^{\text{M}}$. We can conjecture that the following holds:
$\min_{\omega}\mathbb{E}_{(s,a)\sim\mathcal{D}^{\text{M}}}\!\left[-\log\pi^{\text{U}}_{\omega}(a\mid s)\right]\geq\ \min_{\xi}\mathbb{E}_{(s,a)\sim\mathcal{D}^{\text{M}}}\!\left[-\log\pi^{\text{M}}_{\xi}(a\mid s)\right]$
$\Rightarrow D_{\mathrm{KL}}(\beta\|\pi^{\text{U}}_{\omega^{*}})\geq\ D_{\mathrm{KL}}(\beta\|\pi^{\text{M}}_{\xi^{*}})$
As such, we can better effectuate lower OOD rate via Eq. 11 by adopting more expressive policies such as diffusion and flow-matching. It should be noted that the discussions here are agnostic to whether a risk-aware or risk-neutral formulation is adopted.
@@ 4.3 Example
We design a 2‑D contextual bandit with two disjoint modes (Toy Risky Bandit): Top left in Fig. 3 shows a ground truth that consists of safe center (moderate reward, no catastrophic tail) and a risky ring (higher mean with rare large penalties). The task isolates multimodality and lower-tail hazards. Below we introduce our baselines.
@@ 4.3.1 Expressive but risk‑neutral controls
(i) DiffusionQL (Wang et al., 2023): The actor is a conditional diffusion policy. An action is generated by the reverse path $a=\psi_{\theta}(s,z)$ (Eq. 5) with $z\!\sim\!\mathcal{N}(0,I)$. The BC term is the standard noise‑prediction (score‑matching) loss $L^{\text{BC}}_{\text{diff}}(\theta)=\mathbb{E}_{(s,a)\sim\mathcal{D},\ t,\ \varepsilon\sim\mathcal{N}}\!\bigl[\ \|\,\varepsilon-\varepsilon_{\theta}(s,\alpha_{t}a+\sigma_{t}\varepsilon,t)\|_{2}^{2}\ \bigr]$, and risk‑neutral improvement is applied by maximizing the critic’s expected value through the sampler, $L^{\text{Q}}_{\text{diff}}(\theta)\;=\;-\,\mathbb{E}_{s\sim\mathcal{D},\ z\sim\mathcal{N}}\bigl[\,Q_{\phi}\bigl(s,\psi_{\theta}(s,z)\bigr)\,\bigr]$. We use the combined objective $\mathcal{\min}_{\theta}\ L^{\text{BC}}_{\text{diff}}+\eta\,\mathcal{L}^{\text{Q}}_{\text{diff}}$.
(ii) FlowQL (Park et al., 2025): we maintain two policies. The BC flow policy $\mu_{\theta}(s,z)$ is trained only with the flow-matching BC loss (Eq. 6). Alongside, we train a one-step policy $\mu_{\omega}(s,z)$ with the actor loss $L_{\pi}^{\text{FQL}}(\omega)=\mathcal{L}^{\text{distill}}_{\text{flow}}+\eta\,\mathcal{L}^{\text{Q}}_{\text{flow}}$, and deploy $\mu_{\omega}$ at test time (no iterative flow needed).
(iii) Conditional VAE(CVAE)-QL: An autoregressive conditional decoder parameterizes $p_{\theta}(a\mid s)=\prod_{i=1}^{d}p_{\theta}(a_{i}\mid a_{<i},s)$. BC is the NLL $\mathcal{L}^{\text{BC}}_{\text{cvae}}=-\mathbb{E}_{\mathcal{D}}\log p_{\theta}(a\mid s)$, and the risk‑neutral improvement is the same.
@@ 4.3.2 Prior‑anchored perturbation (risk‑aware)
This category of baselines adopt Eq. 12:
(i) ORAAC (Urpí et al., 2021): samples an anchor $b\!\sim\!G_{\phi}(\cdot\mid s)$ from a behavior prior and applies a bounded perturbation toward the actor while optimizing a coherent risk objective (e.g., CVaR);
(ii) ORAAC–Diffusion (Chen et al., 2024): replaces the VAE prior with a diffusion prior, keeping the same anchor–perturb structure;
(iii)ORAAC–Flow: a flow‑prior counterpart in the same mixing form. These (B) baselines let us test the geometric leakage.
@@ 4.3.3 Example Results
Risk‑neutral expressive controls (Fig. 3 b-d): Overall, as expected, these methods are risk-blind and they chase high‑$Q$ ridges without regard to the lower tail. FlowQL often preserves both modes but does not suppress mass on the hazardous ring; Diffusion QL drifts toward sparsely covered high‑$Q$ pockets on the ring, yielding risk exposure; the CVAE variant collapses topology and fills low‑density bridges.
Prior‑anchored perturbation (Fig. 3 e-g): ORAAC and its diffusion/flow variants place substantial mass in the inter‑mode low‑density region, regardless of whether bc prior is expressive or not.
RADAC (Fig. 3 h): by sending CVaR signals from a distributional critic through the diffusion/flow trajectory while regularizing with BC, RADAC concentrates probability near the safe center without filling the gap. Full configuration and additional plots are in App. E.1.
@@
Figure 3: Toy Risky Bandit Results Top: Ground Truth consists of a safe center mode yellow-green and a risky ring where high-reward samples yellow are interspersed with catastrophic penalties (purple). Risk-neutral generative baselines concentrate on the risky ring or collapses topology. Bottom: Prior‑anchored perturbation methods produce samples in the low‑density inter‑mode region, exhibiting OOD leakage. RADAC concentrates near the safe center without losing multimodality. See App. D for more results.
@@ 5 Experiments
In this section, we evaluate RADAC and RAFMAC on the Stochastic-D4RL benchmarks to validate both risk awareness and policy expressiveness. We also quantify the OOD action rate $\varepsilon_{\mathrm{act}}$ (Sec. 5.3) to link practice to our view (Sec 4). Additional results appear in App. D.
@@
We augment standard D4RL locomotion tasks (Fu et al., 2020) with rare heavy‑tailed penalties tied to velocity or torso pitch angles with early termination following  (Urpí et al., 2021) (full construction details and per‑task parameters in App. E). We evaluate on hopper, walker2d, and halfcheetah using the medium–expert and medium–replay datasets, which are multimodal by construction (mixtures of heterogeneous behaviors). It lets us examine that RAMAC learns risk‑aware policies without sacrificing multimodality. During training, we follow  (Urpí et al., 2021) and relabel per-transition rewards in the offline datasets with the same stochastic hazard model; evaluation uses the identical hazard specification.
@@
We compare against representative offline‑RL methods covering value conservatism, distributional conservatism, anchor–perturb risk aversion, and risk–neutral expressive generators: CQL (Kumar et al., 2020), CODAC (Ma et al., 2021), ORAAC (Urpí et al., 2021), DiffusionQL (Wang et al., 2023), and FlowQL (Park et al., 2025).
@@
Following protocols as those adopted in  (Urpí et al., 2021; Wang et al., 2023), we train for 2000 epochs, each with 1000 gradient steps and batch size 256. We evaluate methods at fixed intervals of gradient steps and report (i) raw returns averaged over 5 seeds and (ii) episodic $\mathrm{CVaR}_{0.1}$ computed over 50 rollouts in total (10 evaluation episodes per seed) to avoid normalization bias on the stochastic variants. For ORAAC and CODAC, we adopt the authors’ risk-aware objectives (risk level $\alpha{=}0.1$ unless noted). For the other baselines, we tune hyperparameters within the same training budget to ensure fairness and otherwise use authors’ recommended settings (App. E.3). Further protocol details appear in App. E. For artifact reproducibility, following common practice, we provide full results with corresponding 1000-step evaluation in App. D.
@@ 5.1 Results and Analysis
Table 1 reports Mean and CVaR0.1 for RADAC and RAFMAC alongside baselines. Across six tasks, both RAMAC instantiations deliver strong lower tails with competitive or higher means. Viewed through mode coverage, the diffusion actor in RADAC updates actions over small denoising steps, which tends to maintain diverse in–support modes while letting CVaR guidance move probability away from hazardous boundaries (Dhariwal & Nichol, 2021); this explains RADAC’s stronger CVaR performance on halfcheetah and hopper‑medium‑replay under knife-edge hazards. The flow-matching actor in RAFMAC transports along a short, deterministic ODE path that efficiently sharpens dominant high-reward modes (Lipman et al., 2023) on walker2d‑medium‑expert, often lifting mean return when hazards are smoother. ORAAC regularizes toward a behavior anchor. It reliably handles sharp hazardous thresholds on such as hopper‑medium‑expert but may fail to exploit high-reward modes, and can place mass in low-density between-mode regions depending on the anchor–perturbation. These tendencies align with qualitative safety plots (Fig. 4).
Table 1: Stochastic–D4RL results over 5 seeds. We report Mean and CVaR0.1; best in dark/ second in light shaded. The Full results with s.e. appear in App. D.2.
Dataset Metric CQL CODAC ORAAC FlowQL DiffusionQL RAFMAC RADAC
HalfCheetah-medium-expert Mean 66.66-66.66 0.12-0.12 $796.06$ $844.14$ 20.71-20.71 $889.56$ $\mathbf{916.64}$
CVaR 135.39-135.39 0.11-0.11 $742.94$ $754.44$ 76.39-76.39 $736.95$ $\mathbf{805.25}$
Walker2d-medium-expert Mean 21.52-21.52 $23.96$ $969.62$ $1309.48$ 32.38-32.38 $\mathbf{1822.24}$ $1708.68$
CVaR 64.88-64.88 43.88-43.88 $358.55$ $468.15$ 116.19-116.19 $\mathbf{1127.21}$ $573.22$
Hopper-medium-expert Mean 25.87-25.87 $26.59$ $\mathbf{714.15}$ $341.16$ 279.97-279.97 $281.24$ $130.74$
CVaR 111.37-111.37 150.92-150.92 $\mathbf{374.63}$ 8.80-8.80 872.95-872.95 132.33-132.33 167.29-167.29
HalfCheetah-medium-replay Mean 66.21-66.21 0.11-0.11 $18.99$ $434.33$ $279.95$ $449.04$ $\mathbf{525.84}$
CVaR 127.09-127.09 1.47-1.47 34.09-34.09 $224.73$ $79.93$ $144.73$ $\mathbf{278.65}$
Walker2d-medium-replay Mean 16.90-16.90 $33.59$ $126.94$ $411.36$ $96.88$ 71.69-71.69 $\mathbf{615.94}$
CVaR 51.49-51.49 52.63-52.63 203.64-203.64 $5.08$ $48.14$ $\mathbf{530.37}$ $145.21$
Hopper-medium-replay Mean 16.25-16.25 47.83-47.83 18.00-18.00 $373.16$ 2.79-2.79 $303.44$ $\mathbf{385.58}$
CVaR 118.70-118.70 160.08-160.08 129.25-129.25 62.24-62.24 51.33-51.33 90.73-90.73 $\mathbf{-8.16}$

### v2

@@ 4.1 Prior-Anchored Perturbation and its limitations
where $\zeta_{\psi}$ is a learned residual (optimized to increase $Q$ or CVaR) and the norm bound $\Phi$ keeps updates close to the anchor. Define the anchor support $\mathcal{S}_{G}(s)$ (the region in action space where $G_{\phi}(\cdot\mid s)$ places mass), the full action space $\mathbb{R}^{d}$, and the $\Phi$-radius ball of $b$
@@
where $\mathrm{dist}(x,A)\!\mathrel{\mathop{:}}=\!\inf_{y\in A}\|x-y\|_{2}$ denotes Euclidean distance. OOD can still occur when this margin fails. This method provides a convenient local improvement rule; however, prior work has observed that it suffers from poor mode coverage in multimodal action spaces [wang2022diffusion]. In addition to the identified limitations, we show distinct geometric weakness that can occur even without multimodality; having multiple modes merely magnifies the effect.
Lemma 4.1.
Fix $s$ and write $I_{s}=\mathcal{S}_{G}(s)$ and $O_{s}=\mathbb{R}^{d}\setminus I_{s}$. Suppose there exist an anchor $b^{\star}\in I_{s}$ and a radius $\Phi>0$ such that $\lambda\!\big(B_{\Phi}(b^{\star})\cap O_{s}\big)>0$, and the policy $\pi_{\text{anch}}(\cdot\mid s)$ induced by Eq. 13 admits a density $p(\cdot\mid s)$ with
$p(a\mid s)\;\geq\;c>0\quad\text{for all }a\in B_{\Phi}(b^{\star}).$
Then its per-state OOD probability
$\delta_{s}(\pi_{\text{anch}})\;\mathrel{\mathop{:}}=\;\pi_{\text{anch}}(O_{s}\mid s)$
satisfies
$\delta_{s}(\pi_{\text{anch}})\;\geq\;c\cdot\lambda\bigl(B_{\Phi}(b^{\star})\cap O_{s}\bigr)\;>\;0.$
In particular, as long as the density on $B_{\Phi}(b^{\star})$ remains bounded below by $c>0$, further training of the residual cannot drive $\delta_{s}(\pi_{\text{anch}})$ to zero. Proof appears in App. B.1.
Lemma 4.1 captures how the geometry of the anchor support forces a strictly positive OOD probability once the anchor ball overlaps the complement $O_{s}$. Thin or nonconvex supports make such low-margin anchors $b^{\star}$ unavoidable, and gradients that are not constrained to lie tangent to the data manifold inevitably push some mass across the boundary (see App. B.1.1 for a more detailed discussion).
@@ 4.2 Behavior‑Regularized Objective: Why it works better
This shows that shrinking the forward KL via BC can suppress per–state OOD, with the strength of the effect controlled by $\eta$ (Eq. 12). This helps mitigate the offline RL failure mode discussed in Sec. 2 (proof appears in App. B.2).
We can extend this risk-agnostic mechanism to bound its impact on tail risk. Let $\delta_{s}(\pi_{\theta})$ be the OOD mass at state $s$. Under bounded rewards $Z_{\min}\leq r\leq Z_{\max}$, a simple mixture argument shows that for any risk level $\alpha\in(0,1]$,
$\mathrm{CVaR}_{\alpha}[X_{s}^{\pi}]\;\geq\;\mathrm{CVaR}_{\alpha}[X_{I_{s}}]-\frac{\delta_{s}(\pi_{\theta})}{\alpha}\,(Z_{\max}-Z_{\min}),$
where $X_{I_{s}}$ is the return restricted to the in-support region $I_{s}$. Thus, as long as $\delta_{s}(\pi_{\theta})$ remains small, the worst-case CVaR degradation due to OOD mixing is only linear in the OOD mass. Combined with Proposition 1, this gives a simple CVaR degradation bound controlled by the forward KL in RAMAC.
@@ 4.3 Example
We design a 2‑D contextual bandit with two disjoint modes (Toy Risky Bandit) to make the above geometric analysis concrete: The top-left panel in Fig. 3 shows a ground truth that consists of a safe center (moderate reward, no catastrophic tail) and a risky ring (higher mean with rare large penalties). The task isolates multimodality and lower-tail hazards. Below we introduce our baselines.
@@ 4.3.1 Expressive but risk‑neutral controls
As a common notation, let $G_{\phi}(\cdot\mid s)$ denote an expressive conditional generator (diffusion model, flow, or conditional VAE) that induces the policy $a\sim G_{\phi}(\cdot\mid s)$. Our risk-neutral expressive baselines, DiffusionQL [wang2022diffusion], FlowQL [park2025flow], and a conditional VAE-QL (CVAE-QL), all minimize
$\mathcal{L}_{\mathrm{RN}}(\phi)$ $=\lambda_{\mathrm{BC}}\,\mathbb{E}_{(s,a)\sim\mathcal{D}}\big[\ell\big(a,G_{\phi}(s)\big)\big]-\mathbb{E}_{s\sim\mathcal{D},\,a\sim G_{\phi}(\cdot\mid s)}\big[Q_{\psi}(s,a)\big],$ (15)
i.e., a standard BC loss plus a risk-neutral value-improvement term on top of an expressive actor $G_{\phi}$ with scalar critic $Q_{\psi}$. Each method simply instantiates $G_{\phi}$ (diffusion, flow, or CVAE) and its optimization hyperparameters; full details are given in App. F.1.
@@ 4.3.2 Prior‑anchored perturbation (risk‑aware)
As risk-aware anchor–perturbation baselines we use ORAAC, ORAAC–Diffusion [chen2024uncertainty; urpi2021risk], and ORAAC–Flow, all instantiated via Eq. 13 with their original coherent risk objectives. Concretely, given a behavior anchor $b\sim\beta(\cdot\mid s)$, each method learns a residual perturbation $f_{\phi}$ as in Sec. 4 and optimizes its own risk functional (e.g., CVaR or distorted expectations) under this prior-anchored geometry.
@@ 4.3.3 Example Results
Risk‑neutral expressive controls (Fig. 3b-d): Overall, as expected, these methods are risk-blind and chase high‑$Q$ ridges without regard to the lower tail. FlowQL often preserves both modes but does not suppress mass on the hazardous ring; DiffusionQL drifts toward sparsely covered high‑$Q$ pockets on the ring, yielding risk exposure; the CVAE variant collapses topology and fills low‑density bridges.
Prior‑anchored perturbation (Fig. 3 e-g): ORAAC and its diffusion/flow variants place substantial mass in the inter‑mode low‑density region, regardless of whether the BC prior is expressive or not.
RADAC (Fig. 3 h): By sending CVaR signals from a distributional critic through the diffusion/flow trajectory while regularizing with BC, RADAC concentrates probability near the safe center without filling the gap. Full configuration and additional plots are in App. E.1 and F.1. Overall, these patterns empirically support our theoretical analysis.
@@
Figure 3: Toy Risky Bandit Results Top: Ground truth consists of a safe center mode yellow-green and a risky ring where high-reward samples yellow are interspersed with catastrophic penalties (purple). Risk-neutral generative baselines concentrate on the risky ring or collapse topology. Bottom: Prior‑anchored perturbation methods produce samples in the low‑density inter‑mode region, exhibiting OOD leakage. RADAC concentrates near the safe center without losing multimodality. See App. E.1 for more results.
@@ 5 Experiments
In this section, we evaluate RADAC on the Stochastic-D4RL benchmarks to validate both risk awareness and policy expressiveness. We also quantify the OOD action rate $\varepsilon_{\mathrm{act}}$ (Sec. 5.3) to link practice to the theoretical analysis in Sec. 4. Additional results, including the flow-matching instantiation RAFMAC, appear in App. E.2.
@@
We augment standard D4RL locomotion tasks [fu2020d4rl] with rare heavy‑tailed penalties tied to forward velocity (HalfCheetah) or torso pitch angles (Hopper, Walker2d), together with early termination when the torso leaves a healthy range, following [urpi2021risk]. We evaluate on Hopper, Walker2d, and HalfCheetah using the Medium–Expert and Medium–Replay datasets, which are multimodal by construction. This lets us examine whether RAMAC learns risk‑aware policies without sacrificing multimodality. Full construction details and per‑task parameters appear in App. F.2.
@@
We compare against representative offline‑RL methods covering value conservatism, distributional conservatism, anchor-perturb risk aversion, and risk-neutral expressive generators: CQL [kumar2020conservative], CODAC [ma2021conservative], ORAAC [urpi2021risk], DiffusionQL [wang2022diffusion], and FlowQL [park2025flow]. We defer detailed configurations to App. F.3.
@@
Following the protocols adopted in [urpi2021risk; wang2022diffusion], we train for 2000 epochs, each with 1000 gradient steps. We evaluate methods at fixed intervals of gradient steps and report (i) raw returns averaged over 5 seeds and (ii) episodic $\mathrm{CVaR}_{0.1}$ computed over 50 rollouts in total (10 evaluation episodes per seed) to avoid normalization bias on the stochastic variants. For ORAAC and CODAC, we adopt the authors’ risk-aware objectives (risk level $\alpha{=}0.1$ unless noted). For the other baselines, we tune hyperparameters within the same training budget to ensure fairness and otherwise use authors’ recommended settings. Further protocol details appear in App. F.2. Runtime and inference-latency comparisons for RADAC with DiffusionQL/FlowQL are reported in App. E.4.
@@ 5.1 Results and Analysis
Table 1 reports Mean and CVaR0.1 for RADAC alongside baselines. Across six tasks, RADAC delivers strong lower tails with competitive or higher means. FlowQL is often the strongest risk-neutral baseline, suggesting that flow-based generators can stabilize training even without explicit tail-risk shaping, but its lower-tail returns remain below RADAC. Further analysis, including an ablation on the number of critic quantile particles $N$ and the tail-sample budget $K$ in the CVaR estimator, is reported in App. E.3.2, and a comparison of alternative risk distortions (CVaR vs. Wang and CPW) is given in App. E.3.1. ORAAC regularizes toward a behavior anchor. It reliably handles sharp hazardous thresholds on tasks such as Hopper‑Medium‑Expert but may fail to exploit high-reward modes, and can place mass in low-density regions between modes when the anchor lies in a risky region.
Table 1: Stochastic–D4RL results over 5 seeds. We report Mean and CVaR0.1; best in dark/ second in light shaded. Full results with standard error appear in App. E.2.
Dataset Metric CQL CODAC ORAAC FlowQL DiffusionQL RADAC
HalfCheetah-Medium-Expert Mean 66.66-66.66 0.12-0.12 $796.06$ $844.14$ 20.71-20.71 $\mathbf{916.64}$
CVaR 135.39-135.39 0.11-0.11 $742.94$ $754.44$ 76.39-76.39 $\mathbf{805.25}$
Walker2d-Medium-Expert Mean 21.52-21.52 $23.96$ $969.62$ $1309.48$ 32.38-32.38 $\mathbf{1708.68}$
CVaR 64.88-64.88 43.88-43.88 $358.55$ $468.15$ 116.19-116.19 $\mathbf{573.22}$
Hopper-Medium-Expert Mean 25.87-25.87 $26.59$ $\mathbf{714.15}$ $341.16$ 279.97-279.97 $130.74$
CVaR 111.37-111.37 150.92-150.92 $\mathbf{374.63}$ 8.80-8.80 872.95-872.95 167.29-167.29
HalfCheetah-Medium-Replay Mean 66.21-66.21 0.11-0.11 $18.99$ $434.33$ $279.95$ $\mathbf{525.84}$
CVaR 127.09-127.09 1.47-1.47 34.09-34.09 $224.73$ $79.93$ $\mathbf{278.65}$
Walker2d-Medium-Replay Mean 16.90-16.90 $33.59$ $126.94$ $411.36$ $96.88$ $\mathbf{615.94}$
CVaR 51.49-51.49 52.63-52.63 203.64-203.64 $5.08$ $48.14$ $\mathbf{145.21}$
Hopper-Medium-Replay Mean 16.25-16.25 47.83-47.83 18.00-18.00 $373.16$ 2.79-2.79 $\mathbf{385.58}$
CVaR 118.70-118.70 160.08-160.08 129.25-129.25 62.24-62.24 51.33-51.33 $\mathbf{-8.16}$